# Two-stage LLM Fine-tuning with Less Specialization and More Generalization

### v2

@@ 4 Proposed Method: PROmpt tuning with MOdel Tuning (ProMoT)
ProMoT is a general framework that can be combined with different parameter-efficient tuning and fine-tuning techniques in respective stages. Conceptually, the prompt-tuning at the first stage can be replaced by other commonly used parameter-efficient methods such as LoRA (Hu et al., 2021). However, empirically we found prompt-tuning is much better than LoRA on absorbing format information in early fine-tuning. More discussions can be found in Appendix A.3.6. For fine-tuning methods, we show examples to combine ProMoT with multi-task fine-tuning (Section 5.4) and 1-shot in-context learning prompt (Section 5.3, Section 5.4). Training with 1-shot prompt is introduced by Min et al. (2021) in a multi-task training setting.
@@ 5.1 Settings
We use RTE (Wang et al., 2019; Bentivogli et al., 2009) and WMT14 En-Fr (Bojar et al., 2014) as two fine-tuning tasks in our main experiments. They are selected as examples of classification (RTE) and generative tasks (WMT14 En-Fr translation). Experiments on additional fine-tuning tasks including SNLI (Bowman et al., 2015) and OpenbookQA (Mihaylov et al., 2018) can be found in Appendix A.3.
@@
We primarily use mT5 (Xue et al., 2020) XXL model (Raffel et al., 2020b) in our main experiments, which is pretrained on multi-lingual corpus and contains 13B parameters. This is to accommodate multi-lingual scenarios among our training and evaluation tasks. To show the effectiveness of our method on different pretraining corpus, model sizes and architectures, we also include experiments on mT5 XL, T5.1.1 XXL and PaLM 8b in Appendix A.3. T5 based models are shown to have meaningful few-shot performance as shown in Chung et al. (2022). We do not consider FLAN-T5 (Chung et al., 2022) as a base model in our experiments because it has already been fine-tuned on a large amount of supervised datasets, including our evaluation datasets. More experimental details can be found in Appendix A.2.
@@ 5.2 Supervised Performance on Fine-tuning Tasks
Comparison of supervised performances of a mT5 XXL model on fine-tuning target tasks. We use 0-shot in fine-tuning tasks. We report accuracy for RTE, SNLI and OpenbookQA, and BLEU score for WMT14 En-Fr. We first show that ProMoT training can achieve similar or even better performance on fine-tuning tasks compared to standard fine-tuning. We apply three different fine-tuning methods on four different tasks and report the result in Table 5.2. We report the best performance within the same number of fine-tuning steps (See Appendix A.2 for more details). ProMoT outperforms standard fine-tuning on supervised performance on 3 out of 4 fine-tuning target tasks and outperforms prompt-tuning on 4 out of 4 tasks. Therefore the improved in-context learning performance on unseen tasks (better generalization ability), as will be demonstrated in the next few sections, comes without sacrificing the fine-tune task’s performance.
@@
Performances of a mT5 XXL model finetuned on RTE and evaluated on 8 different tasks to verify the generalization ability. The accuracy on fine-tuned task (RTE) is in the first row. We compare the Norm. Avg. (normalized average performance) with pretrained model and report the relative difference, where red denotes decreased performance and blue denotes increased performance.
5.3 Generalization with Single Task Fine-tuning
@@
In this section, we evaluate and compare the few-shot performance on unseen tasks after fine-tuning. We show the evaluation results of fine-tuning on RTE and WMT14 En-Fr in Table 5.2 and Table 5.3, respectively. Experiments on additional fine-tuning tasks SNLI/OpenbookQA and additional base models including mT5 XL, T5.1.1 XXL and PaLM 8b can be found in Appendix A.3.
@@
[!h] Pretrained Standard Fine-tuning ProMoT (Ours) ProMoT + 1-shot (Ours) Fine-tuning WMT14 En-Fr 1.98 41.80 41.30 41.19 1-shot 4-shots 1-shot 4-shots 1-shot 4-shots 1-shot 4-shots Norm. Avg. 17.52 18.75 9.15 11.67 18.87 20.64 19.91 21.99 (-8.37) (-7.07) (+1.35) (+1.89) (+2.39) (+3.24) CB 46.43 51.79 16.07 32.14 41.07 57.14 41.07 53.57 WiC 49.69 49.69 50.63 49.06 50.16 50.31 49.84 50.63 Evaluation triviaQA 17.58 19.02 3.20 3.15 13.63 15.20 16.93 18.19 web_questions 9.70 13.04 0.89 6.15 9.40 7.92 10.14 12.01 WMT16_ende 3.97 8.83 0.81 0.18 15.52 15.55 16.14 15.63 WMT16_enro 1.82 3.92 1.53 0.42 18.54 17.80 17.57 16.81 XSum 6.41 2.35 0.05 1.86 1.49 0.65 3.41 4.36 WikiLingua/en 4.59 1.33 0.03 0.43 1.14 0.52 4.22 4.73 Performances of a mT5 XXL model finetuned on WMT14 En-Fr and evaluated on 8 few-shot tasks to verify the generalization ability. BLEU on the fine-tuned task is in the first row. We compare the Norm. Avg. (normalized average performance) with pretrained model and report the relative difference, where red denotes decreased performance and blue denotes increased performance.
@@
From both tables, we first observe that the model’s in-context learning performance drops significantly after standard fine-tuning. In particular, the few-shot learning performances drop to near zero for 6 over 8 tasks in Table 5.2, with the only exceptions being CB and WiC where they share the same format (binary classification) as the RTE fine-tuning task.
@@
On the contrary, ProMoT reduces the loss of the in-context learning performance on unseen few-shot evaluation tasks, and even boosts some evaluation tasks that are semantically related to the fine-tuning task but with totally different task formats, resulting in an increasing in-context learning performance on average. In Table 5.2, ProMoT on the binary NLI dataset dataset consistently improves few-shot performances on two summarization tasks beyond the pretrained model. In Table 5.3, ProMoT training on English-French translation substantially improves few-shot performance on other language translation pairs such as English to German and Romanian. This cross-task generalization across different task formats are infeasible with previous fine-tuning techniques. Text examples from standard fine-tuning and ProMoT can be found in Appendix A.3.8. The improvement with less specialization and more generalization can be further boosted when we combine ProMoT with 1-shot prompt to incorporate in-context learning objective during fine-tuning.
@@ 5.5 Ablation Study
6 Conclusions
In this paper, we identify format specialization as one important cause of the loss of general in-context learning abilities during LLM fine-tuning, which tends to happen at the beginning of fine-tuning. We are motivated to develop ProMoT, a simple yet effective two-stage fine-tuning framework that utilizes soft trainable prompts to absorb task-specific formats before model fine-tuning. Experiments on a diverse set of NLP tasks show that ProMoT reduces format specialization and results in surprising generalization across very different tasks, making it a promising method to build general-purpose capabilities into LLMs with small fine-tuning datasets.
Limitations
We have shown the effectiveness of ProMoT in our main paper, however there is no theoretical guarantee on how much format specialization can be absorbed by the soft prompt during the first stage of ProMoT. Besides, our experiments are done with models smaller than 15B due to limited computation resources. It can be interesting to test ProMoT on larger models.
Appendix A Appendix
A.1 Broader Impacts
@@
A.2 Experiment Details
A.2.1 Input Template Used in Experiments
@@
A.2.2 Output Post-processing
@@ A.2.3 Dataset and Models
A.2.4 Hyper-parameters
For all mT5 models, we fine-tune with learning rate 0.001, drop rate 0.1 and label smoothing 0.1, following the default settings for T5 models (Raffel et al., 2020b). For all prompt tuning experiments, we use learning rate 0.2 and prompt length 100. For all tasks except summarization tasks, we choose the model input sequence length larger than the input length in datasets. For summarization, we cut each input to 1024 tokens. We use Adafactor optimizer and batch size 64 without data-packing across all experiments. In inference, we use beam search to decode the outputs with width 4. More experimental settings are provided in the appendix. For ProMoT tuning, at stage 1 we run prompt tuning for 5000 steps and save a checkpoint every 1000 steps, then select the prompt checkpoint with the best performance on target task. At stage 2, we freeze the trained prompt and fine-tune the model for 1000 steps, checkpointing every 100 steps. We pick the model checkpoint with highest performance on the fine-tuned task as our final checkpoint. For comparison, we run prompt tuning and standard fine-tuning for 5000 and 1000 training steps respectively and report the performance of the best checkpoint. We explore fine-tuning with more steps in Appendix A.3.2.
@@ A.2.5 Hardware and Implementation
A.3 Additional Experiment Results
A.3.1 Additional results for single task fine-tuning
@@
A.3.2 Training more steps: trade-off between fine-tuning target task and in-context learning abilities
@@ A.3.4 Additional experiments on mT5 XL
A.3.5 Addtional experiments on PaLM 8B
@@
A.3.6 Using LORA in the first stage
@@
A.3.7 Plotting more steps for Figure 3
@@
A.3.8 Qualitative results on fine-tuning WMT14 En-Fr task

### v3

@@ 4 Proposed Method: PROmpt tuning with MOdel Tuning (ProMoT)
ProMoT is a general framework that can be combined with different parameter-efficient tuning and fine-tuning techniques in respective stages. Conceptually, the prompt-tuning at the first stage can be replaced by other commonly used parameter-efficient methods such as LoRA Hu et al. (2021). However, empirically we found prompt-tuning is much better than LoRA on absorbing format information in early fine-tuning. More discussions can be found in Appendix C.7. For fine-tuning methods, we show examples to combine ProMoT with multi-task fine-tuning (Section 5.4) and 1-shot in-context learning prompt (Section 5.3, Section 5.4). Training with 1-shot prompt is introduced by Min et al. (2021) in a multi-task training setting.
@@ 5.1 Settings
We use RTE (Wang et al., 2019; Bentivogli et al., 2009) and WMT14 En-Fr (Bojar et al., 2014) as two fine-tuning tasks in our main experiments. They are selected as examples of classification (RTE) and generative tasks (WMT14 En-Fr translation). Experiments on additional fine-tuning tasks including SNLI (Bowman et al., 2015) and OpenbookQA (Mihaylov et al., 2018) can be found in Appendix C.
@@
We primarily use mT5 (Xue et al., 2020) XXL model (Raffel et al., 2020b) in our main experiments, which is pretrained on multi-lingual corpus and contains 13B parameters. This is to accommodate multi-lingual scenarios among our training and evaluation tasks. To show the effectiveness of our method on different pretraining corpus, model sizes and architectures, we also include experiments on mT5 XL, T5.1.1 XXL and PaLM 8b in Appendix C. T5 based models are shown to have meaningful few-shot performance as shown in Chung et al. (2022). We do not consider FLAN-T5 (Chung et al., 2022) as a base model in our experiments because it has already been fine-tuned on a large amount of supervised datasets, including our evaluation datasets. More experimental details can be found in Appendix B.
@@ 5.2 Supervised Performance on Fine-tuning Tasks
We first show that ProMoT training can achieve similar or even better performance on fine-tuning tasks compared to standard fine-tuning. We apply three different fine-tuning methods on four different tasks and report the result in Table 5.2. We report the best performance within the same number of fine-tuning steps (See Appendix B for more details). ProMoT outperforms standard fine-tuning on supervised performance on 3 out of 4 fine-tuning target tasks and outperforms prompt-tuning on 4 out of 4 tasks. Therefore the improved in-context learning performance on unseen tasks (better generalization ability), as will be demonstrated in the next few sections, comes without sacrificing the fine-tune task’s performance.
@@
Comparison of supervised performances of a mT5 XXL model on fine-tuning target tasks. We use 0-shot in fine-tuning tasks. We report accuracy for RTE, SNLI and OpenbookQA, and BLEU score for WMT14 En-Fr.
5.3 Generalization with Single Task Fine-tuning
@@
Performances of a mT5 XXL model finetuned on RTE and evaluated on 8 different tasks to verify the generalization ability. The accuracy on fine-tuned task (RTE) is in the first row. We compare the Norm. Avg. (normalized average performance) with pretrained model and report the relative difference, where red denotes decreased performance and blue denotes increased performance. CB and WiC are also NLI tasks, very similar to RTE.
@@
In this section, we evaluate and compare the few-shot performance on unseen tasks after fine-tuning. We show the evaluation results of fine-tuning on RTE and WMT14 En-Fr in Table 5.3 and Table 5.3, respectively. Experiments on additional fine-tuning tasks SNLI/OpenbookQA and additional base models including mT5 XL, T5.1.1 XXL and PaLM 8b can be found in Appendix C.
@@
[!h] Pretrained Standard Fine-tuning ProMoT (Ours) ProMoT + 1-shot (Ours) Fine-tuning WMT14 En-Fr 1.98 41.80 41.30 41.19 1-shot 4-shots 1-shot 4-shots 1-shot 4-shots 1-shot 4-shots Norm. Avg. 17.52 18.75 9.15 11.67 18.87 20.64 19.91 21.99 (-8.37) (-7.07) (+1.35) (+1.89) (+2.39) (+3.24) CB 46.43 51.79 16.07 32.14 41.07 57.14 41.07 53.57 WiC 49.69 49.69 50.63 49.06 50.16 50.31 49.84 50.63 Evaluation triviaQA 17.58 19.02 3.20 3.15 13.63 15.20 16.93 18.19 web_questions 9.70 13.04 0.89 6.15 9.40 7.92 10.14 12.01 WMT16_ende 3.97 8.83 0.81 0.18 15.52 15.55 16.14 15.63 WMT16_enro 1.82 3.92 1.53 0.42 18.54 17.80 17.57 16.81 XSum 6.41 2.35 0.05 1.86 1.49 0.65 3.41 4.36 WikiLingua/en 4.59 1.33 0.03 0.43 1.14 0.52 4.22 4.73 Performances of a mT5 XXL model finetuned on WMT14 En-Fr and evaluated on 8 few-shot tasks to verify the generalization ability. BLEU on the fine-tuned task is in the first row. We compare the Norm. Avg. (normalized average performance) with pretrained model and report the relative difference, where red denotes decreased performance and blue denotes increased performance. WMT16 En-De and En-Ro are translation tasks with different language pairs from WMT14 En-Fr.
@@
From both tables, we first observe that the model’s in-context learning performance drops significantly after standard fine-tuning. In particular, the few-shot learning performances drop to near zero for 6 over 8 tasks in Table 5.3, with the only exceptions being CB and WiC where they share the same format as the RTE fine-tuning task.
@@
On the contrary, ProMoT reduces the loss of the in-context learning performance on unseen few-shot evaluation tasks, and even boosts some evaluation tasks that are semantically related to the fine-tuning task but with totally different task formats, resulting in an increasing in-context learning performance on average. In Table 5.3, ProMoT on the binary NLI dataset dataset consistently improves few-shot performances on two summarization tasks beyond the pretrained model. In Table 5.3, ProMoT training on English-French translation substantially improves few-shot performance on other language translation pairs such as English to German and Romanian. This cross-task generalization across different task formats are infeasible with previous fine-tuning techniques. Text examples from standard fine-tuning and ProMoT can be found in Appendix C.9. The improvement with less specialization and more generalization can be further boosted when we combine ProMoT with 1-shot prompt to incorporate in-context learning objective during fine-tuning.
@@ 5.5 Ablation Study
6 Conclusions and Limitations
In this paper, we identify format specialization as one important cause of the loss of general in-context learning abilities during LLM fine-tuning, which tends to happen at the beginning of fine-tuning. We are motivated to develop ProMoT, a simple yet effective two-stage fine-tuning framework that utilizes soft trainable prompts to absorb task-specific formats before model fine-tuning. Experiments on a diverse set of NLP tasks show that ProMoT reduces format specialization and results in surprising generalization across very different tasks, making it a promising method to build general-purpose capabilities into LLMs with small fine-tuning datasets. Although we have shown the effectiveness of ProMoT in our main paper, there is no theoretical guarantee on how much format specialization can be absorbed by the soft prompt during the first stage of ProMoT. Besides, our experiments are done with models smaller than 15B due to limited computation resources. It can be interesting to test ProMoT on larger models.
Acknowledgments
We thank the reviewers for their invaluable feedbacks. The work is supported in part by NSF 2008173, 2048280, 2331966, ONR N00014-23-1-2300:P00001, ARL 20230936 and Cisco.
Appendix A Broader Impacts
@@
Appendix B Experiment Details
B.1 Input Template Used in Experiments
@@
B.2 Output Post-processing
@@ B.3 Dataset and Models
B.4 Hyper-parameters
For all mT5 models, we fine-tune with learning rate 0.001, drop rate 0.1 and label smoothing 0.1, following the default settings for T5 models (Raffel et al., 2020b). For all prompt tuning experiments, we use learning rate 0.2 and prompt length 100. For all tasks except summarization tasks, we choose the model input sequence length larger than the input length in datasets. For summarization, we cut each input to 1024 tokens. We use Adafactor optimizer and batch size 64 without data-packing across all experiments. In inference, we use beam search to decode the outputs with width 4. More experimental settings are provided in the appendix. For ProMoT tuning, at stage 1 we run prompt tuning for 5000 steps and save a checkpoint every 1000 steps, then select the prompt checkpoint with the best performance on target task. At stage 2, we freeze the trained prompt and fine-tune the model for 1000 steps, checkpointing every 100 steps. We pick the model checkpoint with highest performance on the fine-tuned task as our final checkpoint. For comparison, we run prompt tuning and standard fine-tuning for 5000 and 1000 training steps respectively and report the performance of the best checkpoint. We explore fine-tuning with more steps in Appendix C.3.
@@ B.5 Hardware and Implementation
Appendix C Additional Experiment Results
C.1 Additional results on single task fine-tuning
@@
C.2 4-shot Evaluation Results of Multi-Task training
As an additional result to Table 2, in Table 8 we provide the comparison between the pretrained model, multi-task standard fine-tuning and multi-task ProMoT.
Pretrained multi-task FT Multi ProMoT Multi-task RTE 47.65 90.25 91.34 Fine-tuning WMT14 En-Fr 1.982 41.34 40.73 Norm. Avg. 18.75 20.81 26.31 (+2.06) (+7.56) CB 51.79 82.14 78.57 WiC 49.69 52.82 52.50 Evaluation TriviaQA 19.02 17.75 22.26 Web_questions 13.04 12.25 12.50 WMT16 En-De 8.83 0.24 18.84 WMT16 En-Ro 3.92 0.54 18.81 XSum 2.35 0.34 5.04 WikiLingua/en 1.33 0.39 1.92
Table 8: Comparison of multi-task training on a mixed dataset of RTE and WMT14 En-Fr. We compare the 4-shot evaluation results of pretrained mT5 model, standard multi-task fine-tuning (FT) and multitask (Multi) ProMoT training. We compare the Norm. Avg. (normalized average performance) with pretrained model and report the relative difference, where blue denotes increased performance.
C.3 Training more steps: trade-off between fine-tuning target task and in-context learning abilities
@@ C.5 Additional experiments on mT5 XL
C.6 Addtional experiments on PaLM 8B
@@
C.7 Using LORA in the first stage
@@
C.8 Plotting more steps for Figure 3
@@
C.9 Qualitative results on fine-tuning WMT14 En-Fr task